OpenStudy (rebeccapink6398):

Justin has 6 bags of dog food. His dogs eat 3/4 bag each day. How long will his dog food last?

OpenStudy (anonymous):



OpenStudy (rebeccapink6398):

its 5 1/4

OpenStudy (rebeccapink6398):

I used my iphone 5 siri

OpenStudy (anonymous):

you can also change the denominator

OpenStudy (rebeccapink6398):

>>>???<<<

OpenStudy (rebeccapink6398):

I don't get it

OpenStudy (rebeccapink6398):

@31356 please help :(

OpenStudy (rebeccapink6398):

@jdoe0001 plz help...OH UR HERE!! :D

OpenStudy (jdoe0001):

\(\large { \begin{array}{ccllll} \textit{bags quantity}&days\\ \hline\\ \frac{3}{4}&1\\ 6&x \end{array}\implies \cfrac{\frac{3}{4}}{6}=\cfrac{1}{x} }\) solve for "x"

OpenStudy (jdoe0001):

\(\bf x=\cfrac{6}{\frac{3}{4}}\implies \cfrac{\frac{6}{1}}{\frac{3}{4}}\qquad recall\implies \cfrac{\quad \frac{a}{b}\quad }{\frac{c}{d}}\implies \cfrac{a}{b}\cdot \cfrac{d}{c}\)

OpenStudy (rebeccapink6398):

I don't get it

OpenStudy (jdoe0001):

\(\bf x=\cfrac{6}{\frac{3}{4}}\implies \cfrac{\frac{6}{1}}{\frac{3}{4}}\qquad recall\implies \cfrac{\quad \frac{a}{b}\quad }{\frac{c}{d}}\implies \cfrac{a}{b}\cdot \cfrac{d}{c} \\ \quad \\ thus\implies x=\cfrac{6}{\frac{3}{4}}\implies \cfrac{\frac{6}{1}}{\frac{3}{4}}\implies \cfrac{6}{1}\cdot \cfrac{4}{3}\implies ?\)

OpenStudy (31356):

Let me see......

OpenStudy (31356):

6/(3/4) = 6 * (4/3) = 2*4 = 8?. As a result it is 8 days :D

OpenStudy (31356):

@rebeccapink6398 There's the answer and the steps. Hope that helps! Sorry for answering it too late o.0

OpenStudy (31356):

Medal? If helped?

OpenStudy (31356):

Jdoe is right! :D

OpenStudy (rebeccapink6398):

sorry! I just got done with dinner...And no not too late :)

OpenStudy (rebeccapink6398):

OMG! I just went back (I didn't submit yet) and I put 7! THXXXX @31356 !!!!!

OpenStudy (31356):

My pleasure! :D